# for Hand Motion Generation in Piano Performance

### v1

@@ 4 Baselines
To tackle the challenging task of generating hand motions synchronized with piano audio tracks, we introduce a novel motion generation framework upon the PianoMotion10M dataset. As illustrated in Fig. 3, our framework consists of a position predictor and a gesture generator. The position predictor extracts the hand positions from piano music and integrates them as contextual input for the gesture generator. By leveraging a DDPM model [20], the gesture generator estimates hand pose sequence based on the piano audio and the predicted positions. Further details on each component are elaborated in the following subsections.
@@
Problem Formulation. Given a piano audio piece $A$ with $N$ frames, our objective is to obtain the hand position sequence $P\in\mathbb{R}^{N\times 3\times 2}$ and hand gestures $\Theta\in\mathbb{R}^{N\times J\times 3\times 2}$. The hand position sequence $P$ encompasses the 3D coordinates of the left and right hands. The hand gestures $\Theta$ consist of Euler angle at each joint of both hands.
Due to the highly nonlinear relationship between acoustic signals and hand gestures, it poses a significant challenge to estimate motions through discriminative models [44], which usually leads to an average pose, as demonstrated in Section 5.2. To address this issue, a hand position predictor is introduced to estimate the continuous changes in hand positions. Moreover, a generative model is utilized to reconstruct hand gestures from a piano music piece. Hereby, the task of hand motion generation becomes more concise and comprehensive by disentangling it into hand position estimation and gesture generation. To extract the audio features $f_{a}$ from the audio $A$, we make use of a pre-trained audio feature extractor $\Phi_{a}$ [4, 22], which is formulated as $f_{a}=\Phi_{a}(A),f_{a}\in\mathbb{R}^{N\times C}$. $C$ is the feature dimension of $\Phi_{a}$.
@@ 4.1 Position Predictor
The position predictor is employed to predict the 6-dimensional 3D positions for both the left and right hands. Since the audio features $f_{a}$ cannot directly be mapped to the positions, a feature embedding module is treated as our position decoder $\Phi_{p}$ to extract the latent features. It projects the sequential audio features $f_{a}$ onto the latent features $f_{p}$ with more comprehensive temporal information, which is formulated as $f_{p}=\Phi_{p}(f_{a}),f_{p}\in\mathbb{R}^{N\times 512}$. Subsequently, a linear mapping is employed to project the feature $f_{p}$ onto the output positions $P$.
In our experiments, we can make use of either Transformer [63] or State-Space Model (SSM) [18] as our feature embedding module. Transformer leverages self-attention mechanisms to effectively capture long-range dependencies and contextual information in order to learn temporal relationships. Recently, a different representation inspired by classical SSM [26] is proposed to replace the attention mechanism, which is built upon a more contemporary Selective Structured State Space Model (S6) [18] suitable for deep learning. By sharing a similar architecture to the classical RNN, it can efficiently capture information from previous inputs.
@@ 4.2 Position-guided Gesture Generator
Leveraging recent achievements [76, 3] in motion generation, our approach incorporates a diffusion probabilistic model [20]. By mastering denoising, this model effectively captures the complex distribution of hand motions observed in large-scale piano-motion datasets, which has the capability to generate motions with different conditions. Starting with a clean sample $\Theta_{0}$ of motion sequence, the forward diffusion process establishes a Markov chain that gradually adds noise to $\Theta_{0}$ in $T$ steps, which generates a series of noisy samples $\Theta_{1},...,\Theta_{T}$ as
@@
where $\mathcal{N}$ denotes the Gaussian distribution. $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta$ represent the variance scheduler for the added noise. Therefore, a model parameterized by a deep neural network $G$ is trained to master the inverse process within another Markov chain, which learns the mapping $p(\Theta_{t-1}|\Theta_{t})$ to sequentially denoise samples over $T$ steps. Specifically, denoising model $G$ consists of a 4-layer U-Net [48] with 256, 512, 1024, 2048 dimensions for each layer.
@@ 4.3 Implementation Details
We employ a two-stage scheme to train our proposed network. At the first stage, the position predictor is trained using position error $\mathcal{L}_{p}$ and velocity loss $\mathcal{L}_{v}$. The position error $\mathcal{L}_{p}$ computes the $L_{1}$ loss between the predicted position $\hat{P}$ and the ground truth $P$, expressed as $\mathcal{L}_{p}=||\hat{P}-P||_{1}$. Inspired by [44], velocity loss is adopted to induce temporal stability for generating smoothing movements, which is formulated as $\mathcal{L}_{v}=||(\hat{P}_{n}-\hat{P}_{n-1})-(P_{n}-P_{n-1})||_{2}$. $n$ denotes the $n$-th frame in a motion sequence. Specifically, our model is trained by subject 1467634 and subject 66685747, which have the similar piano keyboard layout. At the second stage, the parameters of position predictor are frozen, and the estimated positions are employed as a guidance for the gesture generator. As in [50], the denoising process is modified from noise prediction to velocity prediction. During the whole training process, the parameters of audio feature extractor $\Phi_{a}$ are frozen.
The baseline model is implemented with PyTorch. We normalize the inputs into 30 FPS through interpolation, where each piece of music lasts 8 seconds. Both stages involve training for $100,000$ iterations at the learning rates of $2e^{-5}$ and $5e^{-5}$, respectively. We conducted all the experiments on a PC with single NVIDIA RTX 3090Ti GPU, which has 24GB of GPU RAM.
@@ 5.1 Evaluation Metrics
To assess the performance of our proposed baseline, we employ several evaluation metrics to examine the effectiveness of hand poses generated by the input piano music, which are crucial in understanding the relationship between piano melody and its corresponding playing motions.
Frechet Inception Distance (FID). Frechet Inception Distance is introduced to measure the Frechet distance between the feature vectors of prediction and ground truth. We pre-train an auto-encoder [70] to project motion sequence onto a latent space for double hands. The FID is adopted to assess the fidelity of the overall motions generated by our baseline model.
Frechet Gesture Distance (FGD). Unlike FID, Frechet Gesture Distance is utilized to compute the disparity between predicted gestures and ground truth of one hand. This metric is instrumental in evaluating the similarity of single hand gestures.
Wasserstein Gesture Distance (WGD). Wasserstein Gesture Distance [49] is computed between two distributions, each of which is represented as a Gaussian Mixture Model (GMM) [27] as below
@@
where $\Pi(\mathcal{I}_{x},\mathcal{I}_{y})$ indicates the joint distributions that combine the parametric GMM distributions $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$. Distributions $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$ are fitted by predicted gestures and ground truth of single hand. $\mathbb{E}$ calculates the expectation of the sample pairs $(x,y)$. The WGD metric offers a robust measure of dissimilarity between the generated motion and ground truth.
Position Distance (PD). The Position Distance calculates the $L_{2}$ distance between estimated positions and the ground truth. It is essential to evaluate the precision of predicted hand positions to ensure the accuracy required for piano fingerings.
@@ 5.2 Experiments
EmoTalk [44] TF - 0.360 0.259 0.033 0.313 0.445 0.232 0.044 0.353 4.645 308
LivelySpeaker [76] TF 1000 0.535 0.249 0.030 0.334 0.538 0.220 0.038 0.406 4.157 321
Our-Base Wav2vec SSM 1000 0.416 0.246 0.034 0.335 0.425 0.223 0.042 0.412 3.587 320
TF 1000 0.424 0.246 0.033 0.334 0.426 0.219 0.040 0.402 3.608 323
HuBERT SSM 1000 0.402 0.247 0.033 0.336 0.432 0.218 0.041 0.407 3.412 320
TF 1000 0.418 0.247 0.034 0.338 0.432 0.219 0.041 0.412 3.529 323
Our-Large Wav2vec SSM 1000 0.421 0.244 0.031 0.208 0.430 0.219 0.038 0.253 3.453 539
TF 1000 0.354 0.244 0.030 0.209 0.372 0.214 0.038 0.251 3.376 557
HuBERT SSM 1000 0.403 0.244 0.030 0.214 0.406 0.217 0.037 0.237 3.395 539
TF 1000 0.351 0.244 0.030 0.205 0.372 0.217 0.037 0.248 3.281 557
Experimental Setup. Within the network, we compare two transformer-based audio feature extractors, Wav2Vec2.0 [4] and HuBERT [22], which are popular in the field of self-supervised speech recognition. Regarding the feature embedding module (Decoder), we evaluate the performance of SSM-based model [18] in contrast to the transformer (TF) approach [63]. We conduct experiments using two different model configurations. Both of them employ the same diffusion architecture while differing in the feature extractor setup. The base model incorporates Wav2Vec2.0-base/HuBERT-base as the audio feature extractors, featuring a model dimension of 768 and containing 8-layer TF/SSM decoder. The large model utilizes Wav2Vec2.0-large/HuBERT-large, with a model dimension of 1,024 and 16-layer TF/SSM decoder.
@@
Quantitative Results. Tab. 3 provides the results of existing methods and our baseline under various experimental settings on PianoMotion10M dataset. Due to the absence of prior work on generating gestures for piano music, we refer to the network structures of EmoTalk [44] and LivelySpeaker [76] and re-implement them to account for our task. EmoTalk directly generates poses from audio, while LivelySpeaker utilizes an MLP-based diffusion backbone for gesture generation. Our baseline models achieve better fidelity on hand motions by estimating positions and gestures, separately. In our baseline models, TF-based models outperform the SSM-based models in processing our time-series information, especially in our large model. For the audio feature extractor, the performance of the HuBERT model slightly outperforms the Wav2vec2. Additionally, we conduct ablation experiments on different denoising steps and our model also achieves competitive results with fewer steps, as shown in Table 4.
@@ Appendix A More Statistics of PianoMotion10M
Figure 5: Distribution of Note Clicks and Volume Levels in the PianoMotion10M Dataset. The top figure depicts note click frequency, and the bottom one shows the volume distribution.
Fig. 5 presents a detailed statistical analysis of piano fingerings, which focuses on the frequency of note clicks and the distribution of volume levels.
Note Click Counts. The top figure in Fig. 5 displays the frequency of each note played, measured in millions. This distribution spans 128 keys of the piano, which indicates frequent usage of those keys in performances. It shows higher counts in specific note ranges.
@@ Appendix B More Visual Samples of PianoMotion10M Dataset
Figure 6: More samples from PianoMotion10M dataset. BV*** denote the corresponding video iDs.
@@ Appendix C More Details of Our Baseline
The audio feature extractor $\Phi_{a}$ maps the audio $A$ to the feature vector $f_{a}$. We use pre-trained Wav2Vec2.0 [4] and HuBERT [22] models developed by Facebook AI for $\Phi_{a}$. Both models leverage extensive unlabeled data for unsupervised pretraining to learn high-dimensional speech representations. HuBERT extends its semi-supervised learning approach with pseudo-labels to self-supervised learning. In our experiments, we use wav2vec2-base-960h and hubert-base-ls960 for the base model, and wav2vec2-large-960h-lv60-self and hubert-large-ls960-ft for the large model as the audio feature extractor $\Phi_{a}$.

### v2

@@ 4 Baselines
To tackle the challenging task of generating hand motions synchronized with piano audio tracks, we introduce a novel motion generation framework upon the PianoMotion10M dataset. As illustrated in Fig. 3, our framework consists of a position predictor and a gesture generator. The position predictor extracts the hand positions from piano music and integrates them as contextual input for the gesture generator. By leveraging a DDPM model (Ho et al., 2020), the gesture generator estimates hand pose sequence based on the piano audio and the predicted positions. Further details on each component are elaborated in the following subsections.
@@
Problem Formulation. Given a piano audio piece $A$ with $N$ frames, our objective is to obtain the hand position sequence $P\in\mathbb{R}^{N\times 3}$ and hand gestures $\Theta\in\mathbb{R}^{N\times J\times 3}$ for both left and right hands.. The hand position sequence $P$ encompasses the 3D coordinates of the left and right hands. The hand gestures $\Theta$ consist of Euler angle at each joint of both hands.
Due to the highly nonlinear relationship between acoustic signals and hand gestures, it poses a significant challenge to estimate motions through discriminative models (Peng et al., 2023), which usually leads to an average pose, as demonstrated in Section 5.2. To address this issue, a hand position predictor is introduced to estimate the continuous changes in hand positions. Moreover, a generative model is utilized to reconstruct hand gestures from a piano music piece. Hereby, the task of hand motion generation becomes more concise and comprehensive by disentangling it into hand position estimation and gesture generation. To extract the audio features $f_{a}$ from the audio $A$, we make use of a pre-trained audio feature extractor $\Phi_{a}$ (Baevski et al., 2020; Hsu et al., 2021), which is formulated as $f_{a}=\Phi_{a}(A),f_{a}\in\mathbb{R}^{N\times C}$. $C$ is the feature dimension of $\Phi_{a}$.
@@ 4.1 Position Predictor
The position predictor is employed to predict the 3D positions of the left and right hands, respectively. Since the audio features $f_{a}$ cannot directly be mapped to the positions, a feature embedding module is treated as our position decoder $\Phi_{p}$ to extract the latent features. It projects the sequential audio features $f_{a}$ onto the latent features $f_{p}$ with more comprehensive temporal information, which is formulated as $f_{p}=\Phi_{p}(f_{a}),f_{p}\in\mathbb{R}^{N\times 512}$. Subsequently, a linear mapping is employed to project the feature $f_{p}$ onto the output positions $P$.
In our experiments, we can make use of either Transformer (Vaswani et al., 2017) or State-Space Model (SSM) (Gu & Dao, 2023) as our feature embedding module. Transformer leverages self-attention mechanisms to effectively capture long-range dependencies and contextual information in order to learn temporal relationships. Recently, a different representation inspired by classical SSM (Kalman, 1960) is proposed to replace the attention mechanism, which is built upon a more contemporary Selective Structured State Space Model (S6) (Gu & Dao, 2023) suitable for deep learning. By sharing a similar architecture to the classical RNN, it can efficiently capture information from previous inputs.
@@ 4.2 Position-guided Gesture Generator
Leveraging recent achievements (Zhi et al., 2023; Azadi et al., 2023) in motion generation, our approach incorporates a diffusion probabilistic model (Ho et al., 2020). By mastering denoising, this model effectively captures the complex distribution of hand motions observed in large-scale piano-motion datasets, which has the capability to generate motions with different conditions. Starting with a clean sample $\Theta_{0}$ of motion sequence, the forward diffusion process establishes a Markov chain that gradually adds noise to $\Theta_{0}$ in $T$ steps, which generates a series of noisy samples $\Theta_{1},...,\Theta_{T}$ as
@@
where $\mathcal{N}$ denotes the Gaussian distribution. $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ and $\beta$ represent the variance scheduler for the added noise. Therefore, a model parameterized by a deep neural network $G$ is trained to master the inverse process within another Markov chain, which learns the mapping $p(\Theta_{t-1}|\Theta_{t})$ to sequentially denoise samples over $T$ steps. Specifically, denoising model $G$ consists of a 4-layer U-Net (Ronneberger et al., 2015) with 256, 512, 1024, 2048 dimensions for each layer.
@@ 4.3 Implementation Details
We employ a two-stage scheme to train our proposed network. At the first stage, the position predictor is trained using position error $\mathcal{L}_{p}$ and velocity loss $\mathcal{L}_{v}$. The position error $\mathcal{L}_{p}$ computes the $L_{1}$ loss between the predicted position $\hat{P}$ and the ground truth $P$, expressed as $\mathcal{L}_{p}=||\hat{P}-P||_{1}$. Inspired by (Peng et al., 2023), velocity loss is adopted to induce temporal stability for generating smoothing movements, which is formulated as $\mathcal{L}_{v}=||(\hat{P}_{n}-\hat{P}_{n-1})-(P_{n}-P_{n-1})||_{2}$. $n$ denotes the $n$-th frame in a motion sequence. Specifically, our model is trained by subject 1467634 and subject 66685747, which have the similar piano keyboard layout. At the second stage, the parameters of position predictor are frozen, and the estimated positions are employed as a guidance for the gesture generator. As in (Salimans & Ho, 2022), the denoising process is modified from noise prediction to velocity prediction. During the whole training process, the parameters of audio feature extractor $\Phi_{a}$ are frozen.
The baseline model is implemented with PyTorch. We normalize the inputs into 30 FPS through interpolation, where each piece of music lasts 8 seconds. Both stages involve training for $100,000$ iterations with the learning rates of $2\times 10^{-5}$ and $5\times 10^{-5}$, respectively. We conducted all the experiments on a PC with single NVIDIA RTX 3090Ti GPU, which has 24GB of GPU RAM.
@@ 5.1 Evaluation Metrics
To assess the performance of our proposed baseline, we employ several evaluation metrics to examine the effectiveness of hand poses generated by the input piano music, which are crucial in understanding the relationship between piano melody and its corresponding playing motions. Specifically, to assess the overall motion distribution similarity of the hand, we employ the Fréchet Inception Distance (FID). In piano performance, the left and right hands execute distinct actions and movements. For a detailed evaluation of the disparity of single-hand movements, we utilize the Fréchet Gesture Distance (FGD) and Wasserstein Gesture Distance (WGD). Evaluating positional accuracy is also crucial, for which we introduce the Positional Distance (PD) as a metric. Smoothness is utilized to determine the fluidity of the generated motions.
Fréchet Inception Distance (FID). FID is introduced to measure the overall motion similarity by calculating the Fréchet distance between the feature vectors of prediction and ground truth. We pre-train an auto-encoder (Yoon et al., 2020) to project motion sequence onto a latent space for double hands. The FID is adopted to assess the fidelity of the overall motions generated by our baseline model. The formulation of FID is derived as follows
$FID=||\mu^{\mathbf{f}}_{pred}-\mu^{\mathbf{f}}_{gt}||^{2}+Tr(C^{\mathbf{f}}_{pred}+C^{\mathbf{f}}_{gt}-2*\sqrt{C^{\mathbf{f}}_{pred}*C^{\mathbf{f}}_{gt}}),$ (3)
where $\mathbf{f}$ represents the latent features obtained by the autoencoder. $\mu$ and $C$ denote the mean vector and covariance matrix, respectively. $Tr()$ is the trace operation of a matrix.
Fréchet Gesture Distance (FGD). On the other hand, FGD is utilized to compute the long-term disparity between predicted gestures and ground truth of one hand. It calculates the Fréchet distance over complete sequences of hand motions $\Theta$, offering a metric for the global alignment between generated and real data. The formula is listed as follows
$FGD=||\mu^{\Theta}_{pred}-\mu^{\Theta}_{gt}||^{2}+Tr(C^{\Theta}_{pred}+C^{\Theta}_{gt}-2*\sqrt{C^{\Theta}_{pred}*C^{\Theta}_{gt}}).$ (4)
Wasserstein Gesture Distance (WGD). WGD (Rubner et al., 2000) is computed between two distributions, each of which is represented as a Gaussian Mixture Model (GMM) (Kolouri et al., 2018) as below
@@
where $\Pi(\mathcal{I}_{x},\mathcal{I}_{y})$ indicates the joint distributions that combine the parametric GMM distributions $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$. Distributions $\mathcal{I}_{x}$ and $\mathcal{I}_{y}$ are fitted by predicted gestures and ground truth of single hand. $\mathbb{E}$ calculates the expectation of the sample pairs $(x,y)$ with multiple Gaussians, which focuses on short-term similarity by analyzing mixture distributions, allowing it to capture finer, localized motion patterns within the sequence. The WGD metric offers a robust measure of dissimilarity between the generated motion and ground truth.
Position Distance (PD). The PD calculates the $L_{2}$ distance between estimated positions and the ground truth, which is formulated as
$PD=||P_{pred}-P_{gt}||^{2}_{2},$ (6)
where $P$ is the hand root position. It is essential to evaluate the precision of predicted hand positions to ensure the accuracy required for piano fingerings.
@@ 5.2 Experiments
EmoTalk (2023) TF - 0.445 0.232 0.044 0.353 0.360 0.259 0.033 0.313 4.645 308
LivelySpeaker (2023) TF 1000 0.538 0.220 0.038 0.406 0.535 0.249 0.030 0.334 4.157 321
Our-Base Wav2vec SSM 1000 0.425 0.223 0.042 0.412 0.416 0.246 0.034 0.335 3.587 320
TF 1000 0.426 0.219 0.040 0.402 0.424 0.246 0.033 0.334 3.608 323
HuBERT SSM 1000 0.432 0.218 0.041 0.407 0.402 0.247 0.033 0.336 3.412 320
TF 1000 0.432 0.219 0.041 0.412 0.418 0.247 0.034 0.338 3.529 323
Our-Large Wav2vec SSM 1000 0.430 0.219 0.038 0.253 0.421 0.244 0.031 0.208 3.453 539
TF 1000 0.372 0.214 0.038 0.251 0.354 0.244 0.030 0.209 3.376 557
HuBERT SSM 1000 0.406 0.217 0.037 0.237 0.403 0.244 0.030 0.214 3.395 539
TF 1000 0.372 0.217 0.037 0.248 0.351 0.244 0.030 0.205 3.281 557
Experimental Setup. Within the network, we compare two transformer-based audio feature extractors, Wav2Vec2.0 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021), which are popular in the field of self-supervised speech recognition. Regarding the feature embedding module (Decoder), we evaluate the performance of SSM-based model (Gu & Dao, 2023) in contrast to the transformer (TF) approach (Vaswani et al., 2017). We conduct experiments using two different model configurations. Both of them employ the same diffusion architecture while differing in the feature extractor setup. The base model incorporates Wav2Vec2.0-base/HuBERT-base as the audio feature extractors, featuring a model dimension of 768 and containing 8-layer TF/SSM decoder. The large model utilizes Wav2Vec2.0-large/HuBERT-large, with a model dimension of 1,024 and 16-layer TF/SSM decoder.
Quantitative Results. Tab. 3 provides the results of existing methods and our baseline under various experimental settings on PianoMotion10M dataset. Due to the absence of prior work on generating gestures for piano music, we refer to the network structures of EmoTalk (Peng et al., 2023) and LivelySpeaker (Zhi et al., 2023) and re-implement them to account for our task. EmoTalk directly generates poses from audio, while LivelySpeaker utilizes an MLP-based diffusion backbone for gesture generation. Our baseline models achieve better fidelity on hand motions by estimating positions and gestures, separately. In our baseline models, TF-based models outperform the SSM-based models in processing our time-series information, especially in our large model. For the audio feature extractor, the performance of the HuBERT model slightly outperforms the Wav2vec2.
@@
Additionally, we conduct ablation experiments on different denoising steps and our model also achieves competitive results with fewer steps, as shown in Table 4.
@@ 6 Conclusion
7 Acknowledgments
This work was supported in part by the National Key Research and Development Program of China under Grant (2023YFF0905104) and in part by the National Natural Science Foundation of China under Grants (62376244). It is also supported by the Information Technology Center and State Key Lab of CAD&CG, Zhejiang University. We are grateful to all Bilibili creators who contributed to the dataset construction, especially CIPMusic, for their support.
@@ Appendix A More Statistics of PianoMotion10M
Figure A1: Distribution of Note Clicks and Volume Levels in the PianoMotion10M Dataset. The top figure depicts note click frequency, and the bottom one shows the volume distribution.
Fig. A1 presents a detailed statistical analysis of piano fingerings, which focuses on the frequency of note clicks and the distribution of volume levels.
Note Click Counts. The top figure in Fig. A1 displays the frequency of each note played, measured in millions. This distribution spans 128 keys of the piano, which indicates frequent usage of those keys in performances. It shows higher counts in specific note ranges.
@@ Appendix B More Visual Samples of PianoMotion10M Dataset
We provide more visual samples of PianoMotion10M dataset, as illustrated in Fig. A2.
Figure A2: More samples from PianoMotion10M dataset. BV*** denote the corresponding video iDs.
@@ Appendix C More Details of Our Baseline
The audio feature extractor $\Phi_{a}$ maps the audio $A$ to the feature vector $f_{a}$. We use pre-trained Wav2Vec2.0 (Baevski et al., 2020) and HuBERT (Hsu et al., 2021) models developed by Facebook AI for $\Phi_{a}$. Both models leverage extensive unlabeled data for unsupervised pretraining to learn high-dimensional speech representations. HuBERT extends its semi-supervised learning approach with pseudo-labels to self-supervised learning. In our experiments, we use wav2vec2-base-960h and hubert-base-ls960 for the base model, and wav2vec2-large-960h-lv60-self and hubert-large-ls960-ft for the large model as the audio feature extractor $\Phi_{a}$.
@@ Appendix D Limitation and Future Work
Figure A3: Samples of failure in the generated hand motions when dealing with excessively fast music.
Our dataset is derived from 14 subjects and faces an issue of imbalance, primarily due to the varying levels of performance provided by authorized creators. In extreme performance scenarios, such as rapid, multi-hand, or non-human piano audio, our baseline method may yield incorrect hand transitions due to out-of-distribution challenges, as illustrated in Fig. A3. Additionally, not all videos meet the required data quality standards. In the future, we plan to engage creators on other platforms, such as YouTube and TikTok, to further enhance the diversity of our dataset.